# Four-dimensional gradient Ricci solitons with (half) nonnegative isotropic curvature

### v2

@@ Abstract.
This is a sequel to our paper [24], in which we investigated the geometry of 4-dimensional gradient shrinking Ricci solitons with half positive (nonnegative) isotropic curvature. In this paper, we mainly focus on 4-dimensional gradient steady Ricci solitons with nonnegative isotropic curvature (WPIC) or half nonnegative isotropic curvature (half WPIC). In particular, for $4$D complete ancient solutions with WPIC, we are able to prove the nonnegativity of the Ricci curvature $Rc\geq 0$ and bound the curvature tensor $Rm$ by $|Rm|\leq R$. For 4D gradient steady solitons with WPIC, we obtain a classification result. We also give a partial classification of 4D gradient steady Ricci solitons with half WPIC. Moreover, we obtain a preliminary classification result for 4D complete gradient expanding Ricci solitons with WPIC. Finally, motivated by the recent work [60], we improve our earlier results in [24] on 4D gradient shrinking Ricci solitons with half PIC or half WPIC, and also provide a characterization of complete gradient Kähler-Ricci shrinkers in complex dimension two among 4-dimensional gradient Ricci shrinkers.
@@ 1. Introduction
$4$-dimensional gradient shrinking Ricci solitons with PIC or WPIC have been classified recently by Li, Ni and Wang [59]. In our very recent work [24], we studied 4-dimensional gradient shrinking Ricci solitons with half PIC or half WPIC. In this paper, we shall explore the geometry and classifications of 4-dimensional complete gradient steady (and expanding) Ricci solitons with WPIC or half WPIC.
@@
Let $(M^{4},g(t)),-\infty<t\leq 0,$ be a 4-dimensional complete ancient solution to the Ricci flow with nonnegative isotropic curvature. Then, $g(t)$ has nonnegative Ricci curvature $Rc\geq 0$ and satisfies the curvature estimate
@@
Furthermore, if $g(t)$ has positive isotropic curvature then it has positive Ricci curvature $Rc>0$.
Remark 1.1.
Curvature estimates in the form of $|Rm|\leq CR$ were first derived by Munteanu-Wang [66] for 4-dimensional gradient shrinking Ricci solitons and subsequently obtained for $4$-dimensional gradient steady solitons in [27], [18, Theorem 5.2], [36] and [28]; see also related results in [23] for 4D Ricci shrinkers and [22] for 4D Ricci expanders.
@@
Every $4$-dimensional complete gradient steady soliton $(M^{4},g,f)$ with WPIC has nonnegative Ricci curvature $Rc\geq 0$ and bounded curvature
@@
    $(M^{4},g,f)$ has positive isotropic curvature and positive Ricci curvature, or
@@
    $(M^{4},g,f)$ is isometric to a quotient of either $N^{3}\times\mathbb{R}$, or $\Sigma^{2}\times\Sigma^{2}$, or $\Sigma^{2}\times\mathbb{R}^{2}$. Here, $\Sigma^{2}$ is the cigar soliton, and $N^{3}$ is either the 3D Bryant soliton or a flying wing.
Remark 1.2.
It follows from Deng-Zhu [39, Theorem 1.2] and a result proved independently by Bryant [14] and Chau-Tam [30] that the gradient steady Kähler-Ricci soliton $(M^{4},g,f)$ in Theorem 1.2(ii) is necessarily biholomorphic to $\mathbb{C}^{2}$.
Since the cigar soliton and flying wings are collapsed, and also there is no $\kappa$-noncollapsed gradient steady Kähler-Ricci soliton with positive holomorphic bisectional curvature by a result of Deng-Zhu [43], we immediately have the following
@@
    $(M^{4},g,f)$ has positive isotropic curvature and positive Ricci curvature, or
@@
Remark 1.3.
By [62, Corollary 5.2], the steady soliton $(M^{4},g,f)$ in Corollary 1.2(a) also has zero asymptotic volume ratio and infinite asymptotic scalar curvature ratio.
Moreover, by combining with [42, Theorem 1.5], we obtain the following result.
Corollary 1.3.
Let $(M^{4},g,f)$ be a $4$-dimensional complete, $\kappa$-noncollapsed, gradient steady Ricci soliton with positive isotropic curvature. Suppose that the scalar curvature $R$ has linear decay so that
$R(x)\leq\frac{C}{1+r(x)}\quad\mbox{on}\ M$
for some constant $C>0$, where $r(x)=d(x_{0},x)$ is the distance function from a fixed point $x_{0}\in M$. Then $(M^{4},g)$ is asymptotically cylindrical, as defined in [11], and the sectional curvature is positive outside a compact set $K\subset M$.
@@
The paper is organized as follows. In Section 2, we recall some basic facts which will be used in the proofs of Theorems 1.11.5. In Section 3, we study 4D gradient steady Ricci solitons with WPIC or half WPIC and prove Theorem 1.1–Theorem 1.3. In particular, we show the somewhat surprising result that any 4-dimensional ancient solution to the Ricci flow with WPIC must have nonnegative Ricci curvature $Rc\geq 0$ (Theorem 3.1) which is essential for deriving the curvature bound $|Rm|\leq R$. In Section 4, we treat the expanding case and prove Theorem 1.4. In Section 5, we examine the shrinking case again and strengthen the results in our previous paper [24] from which Theorem 1.5 follows. Finally, in Section 6, we prove that the property of 2-nonnegative Ricci curvature is preserved by 4D Ricci flow with WPIC and briefly discuss PIC vs. uniformly PIC for 4D ancient solutions.
Acknowledgements. We are grateful to Professor Richard Hamilton for helpful suggestions. We would also like to thank Dr. Yi Lai for helpful comments and Dr. Pak-Yeung Chan for bringing [62, Corollary 5.2] to our attention. The first author’s research was partially supported by a grant from the Simons Foundation.
@@ 2. Preliminaries
2.1. Nonegative isotropic curvature
@@
for any orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}$. For any Riemannian manifold with WPIC, we have the following basic properties (see, e.g., [8, 64]).
@@
Let $(M^{n},g)$, $n\geq 4$, be a Riemannian manifold with nonnegative isotropic curvature. Then, for any orthonormal tangent frame $\{e_{1},\cdots e_{n}\}$ and mutually distinct $i,j,k,l\in\{1,...,n\}$, we have
@@ 2.2. Curvature decomposition of four-manifolds
Let $(M^{4},g)$ be a 4-dimensional (oriented) Riemannian manifold. For the traceless Ricci tensor $\mathring{Rc}$, by choosing a suitable basis for $\wedge^{+}(M)$ and a suitable basis for $\wedge^{-}(M)$, we have
$\begin{split}\lambda_{1}=B_{1}-B_{2}-B_{3},\\ \lambda_{2}=B_{2}-B_{1}-B_{3},\\ \end{split}\quad\quad\begin{split}\lambda_{3}=B_{3}-B_{1}-B_{2},\\ \lambda_{4}=B_{1}+B_{2}+B_{3}.\\ \end{split}$
@@ 2.3. Some basic facts about Ricci solitons and ancient Ricci flows
In the special case of a gradient steady Ricci soliton $(M^{n},g,f)$, by [67, Proposition 3.2], either the scalar curvature $R>0$ everywhere or $(M^{n},g,f)$ is Ricci-flat.
@@
Not only the above result of Chen is very useful, but also Chen’s main idea [^3] Motivated in part by Hamilton’s paper [49, p.2] on eternal solutions to the Ricci flow. of the proof is quite significant and has been widely used in the study of complete ancient solutions; see, e.g., [58, 24, 34]. Very recently, Cho and Li [34] have nicely summarized the essential feature of Chen’s argument and formulated it as the following general lemma (see [34, Corollary 2.4]).
@@
(Chen’s Lemma) Let $(M^{n},g(t))_{t\in(-\infty,0]}$ be an n-dimensional complete ancient solution to the Ricci flow. Suppose $u:M\times(-\infty,0]\rightarrow\mathbb{R}$ is a Lipschitz function and satisfies the differential inequality
@@
    If $g(t)$ has nonnegative isotropic curvature, then it necessarily has nonnegative Ricci curvature $Rc\geq 0$. Moreover, it satisfies the curvature estimate
@@
    If $g(t)$ has positive isotropic curvature, then it necessarily has positive Ricci curvature $Rc>0$.
@@
(a) We first prove that $g(t)$ has $Rc\geq 0$. By (2.4) and Lemma 2.2, the least eigenvalue of $Rc$ is given by $\Lambda_{1}=B_{1}-B_{2}-B_{3}+\frac{R}{4}.$ Thus, $Rc\geq 0$ is equivalent to
(3.3) $u:=R+4B_{1}-4(B_{2}+B_{3})\geq 0.$
Now, we consider the ordinary differential inequality satisfied by the Lipschitz function $u$ in the barrier sense. By Lemma 2.4 and $|\mathring{Rc}|^{2}=4|B|^{2}$,
(3.4) $\begin{split}\frac{d}{dt}u&\geq 2|Rc|^{2}+8[(A_{1}+C_{1})B_{1}+2B_{2}B_{3}]\\ &\quad-8[(A_{2}+C_{2})B_{2}+(A_{3}+C_{3})B_{3}+2(B_{2}+B_{3})B_{1}]\\ &=\frac{1}{2}R^{2}+8|B|^{2}+8(A_{1}+C_{1})B_{1}+16B_{2}B_{3}\\ &\quad-8(A_{2}+C_{2})B_{2}-8(A_{3}+C_{3})B_{3}-16B_{1}(B_{2}+B_{3})\\ \end{split}$
On the other hand, by direct computations, we have
(3.5) $\begin{split}\frac{1}{2}u^{2}&=\frac{1}{2}R^{2}+4RB_{1}+8B_{1}^{2}+8(B_{2}+B_{3})^{2}\\ &\quad-4R(B_{2}+B_{3})-16B_{1}(B_{2}+B_{3}).\end{split}$
Combining (3.4) and (3.5), and using
$R=4(A_{1}+A_{2}+A_{3})=4(C_{1}+C_{2}+C_{3}),$
we get
$\begin{split}\frac{d}{dt}u&\geq\frac{1}{2}u^{2}-4RB_{1}+4R(B_{2}+B_{3})+8(A_{1}+C_{1})B_{1}\\ &\quad-8(A_{2}+C_{2})B_{2}-8(A_{3}+C_{3})B_{3}\\ &=\frac{1}{2}u^{2}-8(A_{1}+A_{2}+A_{3})B_{1}-8(C_{1}+C_{2}+C_{3})B_{1}\\ &\quad+8(A_{1}+A_{2}+A_{3})(B_{2}+B_{3})+8(C_{1}+C_{2}+C_{3})(B_{2}+B_{3})\\ &\quad+8(A_{1}+C_{1})B_{1}-8(A_{2}+C_{2})B_{2}-8(A_{3}+C_{3})B_{3}\\ &=\frac{1}{2}u^{2}+8(B_{2}-B_{1})(A_{3}+C_{3})+8(B_{3}-B_{1})(A_{2}+C_{2})\\ &\quad+8(B_{2}+B_{3})(A_{1}+C_{1})\\ &\geq\frac{1}{2}u^{2}.\end{split}$
Here, in the last inequality, we have used the crucial fact that $A\geq 0$ and $C\geq 0$ from Lemma 2.7(a). Hence, by Lemma 2.6, we conclude that $R+4B_{1}-4(B_{2}+B_{3})\geq 0$, i.e., $Rc\geq 0$.
Next, we prove the curvature estimate $|Rm|\leq R$. By Lemma 2.7, we have $A\geq 0$ and $C\geq 0$. Also, we know that
$\text{tr}\ \!A=\text{tr}\ \!C=R/4.$
Hence, $|A|^{2}\leq\text{tr}(A)^{2}=\frac{1}{16}{R^{2}}$. Similarly, we have $|C|^{2}\leq\frac{1}{16}{R^{2}}$. Next, we know that the Ricci tensor of $(M^{4},g,f)$ is nonnegative. It follows immediately that $|Rc|^{2}\leq R^{2}$. On the other hand,
@@
(3.6) $\begin{split}|Rm|^{2}&\leq 2(|A|^{2}+|C|^{2}+|B|^{2})\\ &\leq\frac{1}{4}R^{2}+2|B|^{2}\leq R^{2}.\end{split}$
This completes the proof of curvature bound $|Rm|\leq R$.
(b) Suppose that $g(t)$ has PIC. We prove $Rc>0$ by contradiction. Assume that $Rc$ has a null eigenvector at some point $(x_{0},t_{0})$. Then, by a standard strong maximum principle argument (cf. [49]) and by observing that if $\Lambda_{1}=0$ attains its minimum at $(x_{0},t_{0})$, then we must have
(3.7) $08(B_{2}-B_{1})(A_{3}+C_{3})+8(B_{3}-B_{1})(A_{2}+C_{2})+8(B_{2}+B_{3})(A_{1}+C_{1})=$ 0
at $(x_{0},t_{0})$. On the other hand, by Lemma 2.7 (b), we know that $A>0$ and $C>0$. Thus, by (3.7), we get
$B_{1}=B_{2}=B_{3}=0.$
However, by (2.4), this would imply that, at $(x_{0},t_{0})$,
$\Lambda_{1}=\frac{R}{4}=A_{1}+A_{2}+A_{3}>0,$
which is a contradiction. Hence $Rc>0$ on $M^{4}\times(-\infty,0]$.
@@
For $n\geq 5$, Li and Ni [58] showed that complete ancient solutions to the Ricci flow with WPIC must have 2-nonnegative Ricci curvature; see [58, Proposition 5.2].
Next, we restate and prove Theorem 1.2.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional complete noncompact, non-flat, gradient steady Ricci soliton with nonnegative isotropic curvature. Then, $(M^{4},g,f)$ either
@@
    $(M^{4},g,f)$ has positive isotropic curvature and positive Ricci curvature, or
@@
    $(M^{4},g,f)$ is a gradient steady Kähler-Ricci soliton with positive holomorphic bisectional curvature, or
@@
    $(M^{4},g,f)$ is isometric to a quotient of either $N^{3}\times\mathbb{R}$, or $\Sigma^{2}\times\Sigma^{2}$, or $\Sigma^{2}\times\mathbb{R}^{2}$. Here, $\Sigma^{2}$ is the cigar soliton and $N^{3}$ is either the 3D Bryant soliton or a flying wing.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional complete noncompact, non-flat, gradient steady Ricci soliton with nonnegative isotropic curvature. Then, by Lemma 2.7, we have $A\geq 0$ and $C\geq 0$.
First of all, we observe that $\ker(A)$ is invariant under parallel translation. Indeed, consider the canonical Ricci flow $g(t)=\Phi(t)^{\ast}(g)$ induced by the steady soliton $(M^{n},g,f)$, with $g(0)=g$, for $t\in[0,1]$. Here, $\Phi(t)$ is the one-parameter family of diffeomorphisms generated by the vector field $\nabla f$. Note that, by Lemma 2.3, the evolution equation of the matrix $A(t)$ for $g(t)$ is given by

### v3

@@ Abstract.
This is a sequel to our paper [24], in which we investigated the geometry of 4-dimensional gradient shrinking Ricci solitons with half positive (nonnegative) isotropic curvature. In this paper, we mainly focus on 4-dimensional gradient steady Ricci solitons with nonnegative isotropic curvature (WPIC) or half nonnegative isotropic curvature (half WPIC). In particular, for $4$D complete ancient solutions with WPIC, we are able to prove the 2-nonnegativity of the Ricci curvature and bound the curvature tensor $Rm$ by $|Rm|\leq R$. For 4D gradient steady solitons with WPIC, we obtain a classification result. We also give a partial classification of 4D gradient steady Ricci solitons with half WPIC. Moreover, we obtain a preliminary classification result for 4D complete gradient expanding Ricci solitons with WPIC. Finally, motivated by the recent work [59], we improve our earlier results in [24] on 4D gradient shrinking Ricci solitons with half PIC or half WPIC, and also provide a characterization of complete gradient Kähler-Ricci shrinkers in complex dimension two among 4-dimensional gradient Ricci shrinkers.
@@ 1. Introduction
Hamilton [49] showed that the PIC (or half PIC) condition is preserved by the Ricci flow in dimension four [^1] Later, this was proved in all dimensions $n\geq 5$ by Brendle-Schoen [12] and Nguyen [65] independently, and this property played an essential role in Brendle-Schoen’s proof of the long standing $1/4$-pinching differentiable sphere theorem [12]. and successfully initiated the investigation of 4D Ricci flow with surgery under the PIC assumption; see also the work of Chen-Zhu [32]. Subsequently, by using the Ricci flow with surgery developed in [49, 32], Chen-Tang-Zhu [31] classified compact $4$-manifolds with PIC up to diffeomorphisms.
Meanwhile, $4$-dimensional gradient shrinking Ricci solitons with PIC or WPIC have been classified recently by Li, Ni and Wang [58]. In our very recent work [24], we studied 4-dimensional gradient shrinking Ricci solitons with half PIC or half WPIC. In this paper, we shall explore the geometry and classifications of 4-dimensional complete gradient steady (and expanding) Ricci solitons with WPIC or half WPIC.
@@
Let $(M^{4},g(t)),-\infty<t\leq 0,$ be a 4-dimensional complete ancient solution to the Ricci flow with nonnegative isotropic curvature. Then, $g(t)$ has 2-nonnegative Ricci curvature and satisfies the curvature estimate
@@
Furthermore, if $g(t)$ has positive isotropic curvature then it has 2-positive Ricci curvature.
@@
Every $4$-dimensional complete gradient steady soliton $(M^{4},g,f)$ with WPIC has 2-nonnegative Ricci curvature and bounded curvature
@@
Remark 1.1.
Curvature estimates in the form of $|Rm|\leq CR$ for 4D gradient Ricci solitons were first derived by Munteanu-Wang [64] in the shrinking case (with bounded scalar curvature) and subsequently obtained for $4$D gradient steady solitons in [20], [27], [17, Theorem 5.2], [36] and [28] under certain conditions; see also related results in [23] for 4D Ricci shrinkers and [22] for 4D Ricci expanders.
@@
    $(M^{4},g,f)$ has PIC and 2-positive Ricci curvature, or
@@
    $(M^{4},g,f)$ is isometric to a quotient of either $N^{3}\times\mathbb{R}$, or $\Sigma^{2}\times\Sigma^{2}$, or $\Sigma^{2}\times\mathbb{R}^{2}$. Here, $\Sigma^{2}$ is the cigar soliton; $N^{3}$ is the Bryant soliton or a flying wing.
Since the cigar soliton and 3D flying wings are collapsed, and also there is no $\kappa$-noncollapsed gradient steady Kähler-Ricci soliton with positive holomorphic bisectional curvature by a result of Deng-Zhu [42], we immediately have the following
@@
    $(M^{4},g,f)$ has PIC and 2-positive Ricci curvature, or
@@
The paper is organized as follows. In Section 2, we recall some basic facts which will be used in the proofs of Theorems 1.11.5. In Section 3, we study 4D gradient steady Ricci solitons with WPIC or half WPIC and prove Theorem 1.1–Theorem 1.3. In particular, we show that any 4-dimensional ancient solution to the Ricci flow with WPIC must have 2-nonnegative Ricci curvature (Theorem 3.1(a)) which is essential for deriving the curvature bound $|Rm|\leq R$. In Section 4, we treat the expanding case and prove Theorem 1.4. In Section 5, we examine the shrinking case again and strengthen the results in our previous paper [24] from which Theorem 1.5 follows. Finally, in Section 6, we prove that the property of 2-nonnegative Ricci curvature is preserved by 4D Ricci flow with WPIC and briefly discuss PIC vs. uniformly PIC for 4D ancient solutions.
@@ 2. Preliminaries
2.1. Nonnegative isotropic curvature
@@
for any orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}$. For Riemannian manifolds with WPIC, we have the following basic properties.
@@
(Brendle [8] & Micallef-Wang [62]) Let $(M^{n},g)$, $n\geq 4$, be a Riemannian manifold with nonnegative isotropic curvature. Then, for any orthonormal tangent frame $\{e_{1},\cdots e_{n}\}$ and mutually distinct $i,j,k,l\in\{1,...,n\}$, we have
@@ 2.2. Curvature decomposition of four-manifolds
Let $(M^{4},g)$ be a 4-dimensional (oriented) Riemannian manifold. Then, the sum of the least two eigenvalues of $Rc$ is given by
$\Lambda_{1}+\Lambda_{2}=\frac{1}{2}(R-4B_{3}).$
Proof.
It is well known (see, e.g., equation [24, (2.7)]) that
$\begin{split}\mathring{Rc}=\begin{pmatrix}B_{11}+B_{22}+B_{33}&B_{32}-B_{23}&B_{13}-B_{31}&B_{21}-B_{12}\\ B_{32}-B_{23}&B_{11}-B_{22}-B_{33}&B_{21}+B_{12}&B_{13}+B_{31}\\ B_{13}-B_{31}&B_{21}+B_{12}&B_{22}-B_{11}-B_{33}&B_{23}+B_{32}\\ B_{21}-B_{12}&B_{13}+B_{31}&B_{23}+B_{32}&B_{33}-B_{11}-B_{22}\end{pmatrix},\end{split}$
where $B_{ij}$ are the entries of the matrix $B$ in (2.3). Now, we choose a frame $\{e_{1},e_{2},e_{3},e_{4}\}$ such that the traceless Ricci tensor $\mathring{Rc}$ is diagonoal so that the matrix $B$ is also diagonal, with $B_{11}\leq B_{22}\leq B_{33}$. Then, we have
(2.5) $\begin{split}\lambda_{1}=B_{11}-B_{22}-B_{33},\\ \lambda_{2}=B_{22}-B_{11}-B_{33},\\ \end{split}\quad\quad\begin{split}\lambda_{3}=B_{33}-B_{11}-B_{22},\\ \lambda_{4}=B_{11}+B_{22}+B_{33}.\end{split}$
In particular,
(2.6) $\lambda_{1}+\lambda_{2}=-2B_{33}.$
On the other hand, by equation [24, (2.5)], we have
$\begin{split}4B_{11}=\Lambda_{4}+\Lambda_{1}-\Lambda_{2}-\Lambda_{3},\\ 4B_{22}=\Lambda_{4}+\Lambda_{2}-\Lambda_{3}-\Lambda_{1},\\ 4B_{33}=\Lambda_{4}+\Lambda_{3}-\Lambda_{1}-\Lambda_{2},\\ \end{split}$
which imply $B_{33}\geq B_{22}\geq 0$. Since $0\leq B_{1}\leq B_{2}\leq B_{3}$ are the singular eigenvalues of $B$, i.e., they are the eigenvalues of the matrix $\sqrt{B{}^{t}B}$, it follows that
(2.7) $B_{1}=|B_{11}|=\pm B_{11},\quad B_{2}=|B_{22}|=B_{22},\quad B_{3}=|B_{33}|=B_{33}.$
Therefore, Lemma 2.2 follows immediately from (2.4), (2.6), and (2.7). ∎
@@ 2.3. Some basic facts about Ricci solitons and ancient Ricci flows
In the special case of a gradient steady Ricci soliton $(M^{n},g,f)$, by [66, Proposition 3.2], either the scalar curvature $R>0$ everywhere or $(M^{n},g,f)$ is Ricci-flat.
@@
Not only the above result of Chen is very useful, but also Chen’s main idea [^4] Motivated in part by Hamilton’s paper [48, p.2] on eternal solutions to the Ricci flow. of the proof is quite significant and has been widely used in the study of complete ancient solutions; see, e.g., [57, 24, 34]. Very recently, Cho and Li [34] have nicely summarized the essential feature of Chen’s argument and formulated it as the following general lemma.
@@
(Chen’s Lemma [34, Corollary 2.4]) Let $(M^{n},g(t))_{t\in(-\infty,0]}$ be an n-dimensional complete ancient solution to the Ricci flow. Suppose $u:M\times(-\infty,0]\rightarrow\mathbb{R}$ is a Lipschitz function and satisfies the differential inequality
@@
    If $g(t)$ has nonnegative isotropic curvature, then it has 2-nonnegative Ricci curvature. Moreover, it satisfies the curvature estimate
@@
    If $g(t)$ has positive isotropic curvature, then it has 2-positive Ricci curvature.
@@
(a) We first prove that $g(t)$ has 2-nonnegative Ricci curvature. By Lemma 2.2, the sum of the least two eigenvalues of $Rc$ is given by
$\Lambda_{1}+\Lambda_{2}=\frac{1}{2}(R-4B_{3}).$
Thus, 2-nonnegative Ricci curvature is equivalent to $u:=R-4B_{3}\geq 0.$
Now, we consider the ordinary differential inequality satisfied by the Lipschitz function $u$ in the barrier sense. By Lemma 2.4 and $|\mathring{Rc}|^{2}=4|B|^{2}$, we have
$\begin{split}\frac{d}{dt}u&\geq 2|Rc|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\\ &=\frac{1}{2}R^{2}+8|B|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\\ &=\frac{1}{2}\left(R-4B_{3}\right)^{2}+4RB_{3}+8(B_{2}-B_{1})^{2}-8(A_{3}+C_{3})B_{3}\\ &\geq\frac{1}{2}u^{2}+4(R-2A_{3}-2C_{3})B_{3}.\end{split}$
On the other hand, since $A_{1}+A_{2}+A_{3}=C_{1}+C_{2}+C_{3}=\frac{R}{4}$, we have
$R-2A_{3}-2C_{3}=2(A_{1}+A_{2}+C_{1}+C_{2}).$
Hence, we obtain
(3.3) $\begin{split}\frac{d}{dt}u&\geq\frac{1}{2}u^{2}+8(A_{1}+A_{2}+C_{1}+C_{2})B_{3}\\ &\geq\frac{1}{2}u^{2},\end{split}$
where, in the last inequality, we have used the assumption of WPIC, i.e., $A_{1}+A_{2}\geq 0$ and $C_{1}+C_{2}\geq 0$. Hence, by Lemma 2.6, we have $R-4B_{3}\geq 0$, i.e., $Rc$ is 2-nonnegative.
Next, we prove the curvature estimate $|Rm|\leq R$. By Lemma 2.7, we have $A\geq 0$ and $C\geq 0$. Also, we know that $\text{tr}A=\text{tr}C=R/4.$ Hence, $|A|^{2}\leq\text{tr}(A)^{2}=\frac{1}{16}{R^{2}}$. Similarly, we have $|C|^{2}\leq\frac{1}{16}{R^{2}}$. Since the Ricci tensor of $(M^{4},g,f)$ is 2-nonnegative, it is not hard to see that $|Rc|^{2}\leq R^{2}$. Indeed, let $\Lambda_{1}\leq\Lambda_{2}\leq\Lambda_{3}\leq\Lambda_{4}$ be the eigenvalues of $Rc$, then, as $\Lambda_{1}+\Lambda_{2}\geq 0$, we obtain
$\begin{split}|Rc|^{2}&=\sum_{i=1}^{4}\Lambda^{2}_{i}\leq 2\Lambda_{2}^{2}+\Lambda_{3}^{2}+\Lambda_{4}^{2}\leq(\Lambda_{3}+\Lambda_{4})^{2}\\ &\leq\left(\sum_{i=1}^{4}\Lambda_{i}\right)^{2}=R^{2}.\end{split}$
On the other hand,
@@
(3.4) $\begin{split}|Rm|^{2}&\leq 2(|A|^{2}+|C|^{2}+|B|^{2})\\ &\leq\frac{1}{4}R^{2}+2|B|^{2}\leq\frac{5}{8}R^{2}.\end{split}$
This completes the proof of curvature bound.
(b) Suppose that $g(t)$ has PIC. We prove 2-positive Ricci curvature by contradiction. We consider the quadratic form $Z:=RI-4\sqrt{B{}^{t}B}$, where $I$ is the 3 by 3 identity matrix. By part (a), we know that $Z\geq 0$ and that 2-positive Ricci curvature is equivalent to $Z>0$. Now, we denote the eigenvalues of $Z$ by
(3.5) $0\leq Z_{1}\leq Z_{2}\leq Z_{3}.$
Assume that $Z$ has a null eigenvector at some space-time point $(x_{0},t_{0})$. Then, by using a similar argument as in the proof of [24, Proposition 3.1(b)], we get a contradiction. Alternatively, by a standard strong maximum principle argument (cf. [48]), it follows from (3.3) that if $Z_{1}=0$ attains its minimum at $(x_{0},t_{0})$, we must have
(3.6) $A_{1}+A_{2}+C_{1}+C_{2}=0$
at $(x_{0},t_{0})$. On the other hand, by Lemma 2.7 (b), we know that $A>0$ and $C>0$. This is a contradiction. Hence, the Ricci tensor is 2-positive on $M^{4}\times(-\infty,0]$.
@@
For $n\geq 5$, Li and Ni [57] proved that $n$-dimensional complete ancient solutions to the Ricci flow with WPIC must have 2-nonnegative Ricci curvature; see [57, Proposition 5.2]. However, their proof does not extend to the $n=4$ case, since the key inequality [57, Lemma 4.1] in the proof holds only for $n\geq 5$.
Proof of Theorem 1.2. Let $(M^{4},g,f)$ be a $4$-dimensional complete noncompact, non-flat, gradient steady Ricci soliton with nonnegative isotropic curvature.
First of all, we claim that if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$ then $(M^{4},g,f)$ has PIC. Indeed, consider the canonical Ricci flow $g(t)=\Phi(t)^{\ast}(g)$ induced by the steady soliton $(M^{n},g,f)$, with $g(0)=g$, for $t\in[0,1]$. Here, $\Phi(t)$ is the one-parameter family of diffeomorphisms generated by the vector field $\nabla f$. By [12, Proposition 8], for any fixed time $t\in(0,1)$, the set of all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{4},g(t))$ satisfy
(3.7) $R_{1313}+R_{1414}+R_{2323}+R_{2424}-2R_{1234}=0$
is invariant under parallel transport with respect to the metric $g(t)$. Moreover, as $g(t)=\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{n},g,f)$ satisfying (3.7) is also invariant under parallel translation with respect to $g$. Thus, by [8, Corollary 9.14], if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$ then $(M^{4},g,f)$ has PIC.
Now, we are ready to conclude our proof.
Case 2: $(M^{4},g,f)$ is locally irreducible and locally symmetric. Then, as $(M^{4},g,f)$ is a steady Ricci soliton it must be Ricci flat. Together with the third inequality in Lemma 2.1, or by [5, Theorem 7.61], this would imply $(M^{4},g,f)$ is flat. So, Case 2 cannot occur when $(M^{4},g,f)$ is assumed non-flat.
Case 3: $(M^{4},g,f)$ is locally irreducible and not locally symmetric. Then, by Berger’s holonomy classification, either $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$, or $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, or $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$. If $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$ then, from the above, we know that $(M^{4},g,f)$ must have PIC and 2-positive Ricci curvature. On the other hand, if $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$ then $(M^{4},g,f)$ is Calabi-Yau, hence flat due to WPIC. But this is ruled out by the non-flatness assumption. Finally, if $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, then $(M^{4},g)$ is Kähler, hence $(M^{4},g,f)$ is a gradient steady Kähler-Ricci soliton. On the other hand, any gradient steady Kähler-Ricci soliton of complex dimension two with WPIC necessarily has nonnegative curvature operator; see, e.g., [34, Lemma 4.6]. Moreover, since $(M^{4},g,f)$ is locally irreducible, not locally symmetric and has $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, by using [45, Theorem 8.3] and a similar argument as in [19], it follows that $(M,g)$ has $Rm>0$ when restricted to real $(1,1)$-forms $\wedge^{1,1}_{\mathbb{R}}(M)$. This completes the proof of Theorem 1.2. $\square$
Finally, we restate and prove Theorem 1.3.
@@
    $(M^{4},g,f)$ has half positive isotropic curvature, or
@@
    $(M^{4},g,f)$ is Calabi-Yau, hence locally hyperKähler, or
@@
    $(M^{4},g,f)$ is Ricci-flat and self-dual, or
  • (iv)
  • (v)
@@
First of all, by Lemma 2.7, $(M^{4},g,f)$ has either $A\geq 0$ or $C\geq 0$. Without loss of generality, we may assume $A\geq 0$.
Next, we observe that $\ker(A)$ is invariant under parallel translation. Indeed, as in the proof of Theorem 1.2, we consider the canonical Ricci flow $g(t)=\Phi(t)^{\ast}(g)$ induced by the steady soliton $(M^{n},g,f)$, with $g(0)=g$, for $t\in[0,1]$. Note that, by Lemma 2.3, the evolution equation of the matrix $A(t)$ for $g(t)$ is given by